# IG Captioner: Information Gain Captioners are Strong Zero-shot Classifiers

### v1

@@ 5.5 Weights of the Multimodal and Unimodal Training Losses
Settings. We ablate the multimodal loss weight $\beta$ and the unimodal loss weight $\gamma$ in the training objective of IG captioner (Eqn. (6)). We train the IG captioners with ViT-B encoder on the Laion-5B dataset and evaluate them on zero-shot ImageNet classification.
Results. The results are shown in Tab. 8. The equal weights do not produce the optimal results. We notice that image encoder is only trained with the multimodal loss while the text decoder is supervised by both the losses. We set the multimodal loss weight $\beta=1.5$ and the unimodal loss $\gamma=0.5$ for all the other experiments.
@@ 6 Conclusion
In this paper, we tackle the challenging problem of using a generative captioner to perform zero-shot classification tasks. We identify that the captioner is negatively impacted by linguistic priors and likely to ignore the visual signals. To reduce such impact, we propose the Information Gain (IG) captioner with the novel and effective approaches for both the training and evaluation processes. IG captioner demonstrates to be a strong zero-shot classifier. We hope that this paper will encourage further research on developing more powerful zero-shot generative classifiers.

### v2

@@ 5.5 Weights of the Multimodal and Unimodal Training Losses
models VQAv2 test-std acc. (%)
CLIP N/A
captioner 74.5
IG captioner 75.2
Table 10: Generative evaluation on VQAv2. All the models are pretrained on the Laion-5B dataset with ViT-L encoder and fine-tuned on the VQAv2 dataset. CLIP can not perform the VQA task as a discriminative model.
5.5.1 Settings and Results.
We ablate the multimodal loss weight $\beta$ and the unimodal loss weight $\gamma$ in the training objective of IG captioner (Eqn. (6)). We train the IG captioners with ViT-B encoder on the Laion-5B dataset and evaluate them on zero-shot ImageNet classification. The results are shown in Tab. 10. The equal weights do not produce the optimal results. We notice that image encoder is only trained with the multimodal loss while the text decoder is supervised by both the losses. We set the multimodal loss weight $\beta=1.5$ and the unimodal loss $\gamma=0.5$ for all the other experiments.
5.6 Generative Evaluation of IG Captioner
5.6.1 Settings and Results.
We evaluate the generative capability of IG captioner in this section. Specifically, after pertaining on the Laion-5B dataset, we perform the generative fine-tuning on the VQAv2 dataset. The results are shown in Tab. 10. The IG captioner outperforms the captioner, showing that the information gain approach improves the captioner in both discriminative and generative tasks.
@@ 6 Conclusion
In this paper, we address the challenge of using a generative captioner for zero-shot classification tasks. We identify and mitigate the issue of linguistic priors dominating visual signals by introducing the Information Gain (IG) captioner with novel training and evaluation methods. Our results show that the IG captioner is an effective zero-shot classifier.
6.0.1 Acknowledgements
We gratefully acknowledge supports from ONR with N00014-23-1-2641.